# Diffusion Attribution Score: Evaluating Training Data Influence in Diffusion Model

### v2

@@ Abstract
As diffusion models become increasingly popular, the misuse of copyrighted and private images has emerged as a major concern. One promising solution to mitigate this issue is identifying the contribution of specific training samples in generative models, a process known as data attribution. Existing data attribution methods for diffusion models typically quantify the contribution of a training sample by evaluating the change in diffusion loss when the sample is included or excluded from the training process. However, we argue that the direct usage of diffusion loss cannot represent such a contribution accurately due to the calculation of diffusion loss. Specifically, these approaches measure the divergence between predicted and ground truth distributions, which leads to an indirect comparison between the predicted distributions and cannot represent the variances between model behaviors. To address these issues, we aim to measure the direct comparison between predicted distributions with an attribution score to analyse the training sample importance, which is achieved by Diffusion Attribution Score (DAS). Underpinned by rigorous theoretical analysis, we elucidate the effectiveness of DAS. Additionally, we explore strategies to accelerate DAS calculations, facilitating its application to large-scale diffusion models. Our extensive experiments across various datasets and diffusion models demonstrate that DAS significantly surpasses previous benchmarks in terms of the linear data-modelling score, establishing new state-of-the-art performance.
@@ 1 Introduction
Diffusion models, highlighted in key studies (Ho et al., 2020; Song et al., 2021b), are advancing significantly in generative machine learning with broad applications from image generation to artistic creation (Saharia et al., 2022; Hertz et al., 2023; Li et al., 2022; Ho et al., 2022). As these models, exemplified by projects like Stable Diffusion (Rombach et al., 2022), become increasingly capable of producing high-quality, varied outputs, the misuse of copyrighted and private images has become a significant concern. A key strategy to address this issue is identifying the contributions of training samples in generative models by evaluating their influence on the generated images, a task known as data attribution. Data attribution in machine learning is essential for tracing model outputs back to influential training examples and understanding how specific data points affect model behavior. In practical applications, data attribution spans various domains, including explaining predictions (Koh & Liang, 2017; Yeh et al., 2018; Ilyas et al., 2022), curating datasets (Khanna et al., 2019; Jia et al., 2021; Liu et al., 2021), and dissecting the mechanisms of generative models like GANs and VAEs (Kong & Chaudhuri, 2021; Terashita et al., 2021), serving to enhance model transparency and explore the effect of training data on model behaviors.
@@
Data attribution methods generally fall into two categories. The first, based on sampling (Shapley et al., 1953; Ghorbani & Zou, 2019; Ilyas et al., 2022), involves retraining models to assess how outputs change with the deletion of specific data. While effective, this method requires training thousands of models—a formidable challenge given today’s large-scale models. The second approach uses approximations to assess the change in output for efficiency (Koh & Liang, 2017; Feldman & Zhang, 2020; Pruthi et al., 2020), which may compromise precision. Recent advancements have led to innovative estimators like TRAK (Park et al., 2023), which linearizes model behavior using a kernel matrix analogous to the Fisher Information Matrix. Additionally, Zheng et al. (2024) proposed D-TRAK, which adapts TRAK to diffusion models by setting the output function as the diffusion loss. However, this setting conducts an indirect comparison between the predicted distributions since the diffusion loss is equivalent to the divergence between predicted and ground truth distributions (Ho et al., 2020). They also reported counterintuitive findings that the diffusion loss can be replaced with other output functions without modifying the form of the attribution score, and the results of these empirical designs surpass diffusion loss with theoretical designs, highlighting the need for a deeper understanding of diffusion models in data attribution.
In this paper, we first present a theoretical analysis to address the challenges associated with directly applying TRAK to diffusion models, which also explain the counter-intuitive observations noted in D-TRAK. Further, we propose a novel attribution method termed Diffusion Attribution Score (DAS), specifically designed for diffusion models to quantitatively assess the impact of training samples on model outputs by measuring the divergence between predicted distributions when the sample is included or excluded from the training set, which includes a self-contained derivation. Additionally, we explore several techniques to expedite the computation of DAS, such as compressing models or datasets, enabling its application to large-scale diffusion models and significantly enhancing the practicability of our approach. The primary contributions of our work are summarized as follows:
@@
    We offer a comprehensive analysis of the limitations when directly applying TRAK to diffusion tasks. This theoretical examination sheds light on the empirical success of D-TRAK and fosters the development of more effective attribution methods.
@@ 2.1 Diffusion Models
Our study concentrates on discrete-time diffusion models, specifically Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) and Latent Diffusion Models (LDMs) which are foundational to Stable Diffusion (Rombach et al., 2022). This paper grounds all theoretical derivations related to diffusion models within the framework of unconditional generation using DDPMs. Below, we detail the notation employed in DDPMs that underpins all further theoretical discussions.
Consider a training set ${\mathbb{S}}=\{{\bm{z}}^{(1)},...,{\bm{z}}^{(n)}\}$ where each training sample ${\bm{z}}^{(i)}:=({\bm{x}}^{(i)},y^{(i)})\sim\mathcal{Z}$ is an input-label pair [^1] In text-to-image task, ${\bm{z}}:=({\bm{x}}^{i},y^{i})$ represents an image-caption sample, whereas in unconditional generation, it solely contains an image ${\bm{z}}:=({\bm{x}})$.. Given an input data distribution $q({\bm{x}})$, DDPMs aim to model a distribution $p_{\theta}({\bm{x}})$ to approximate $q({\bm{x}})$. The learning process is divided into forward and reverse process, conducted over a series of timesteps in the latent variable space, with ${\bm{x}}_{0}$ denoting the initial image and ${\bm{x}}_{t}$ the latent variables at timestep $t\in[1,T]$. In the forward process, DDPMs sample an observation ${\bm{x}}_{0}$ from ${\mathbb{S}}$ and progressively add noise on it across $T$ timesteps: $q({\bm{x}}_{t}|{\bm{x}}_{t-1}):=\mathcal{N}({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},\beta_{t}{\bm{I}})$, where $\beta_{1},...,\beta_{T}$ constitute a variance schedule. As indicated in DDPMs, the latent variable ${\bm{x}}_{t}$ can be express as a linear combination of ${\bm{x}}_{0}$:
@@
where $\alpha_{t}:=1-\beta_{t},\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$ and ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$. In the reward process, DDPMs model a distribution $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ by minimizing the KL-divergence from $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$ at each timestep $t$:
@@
where ${\bm{\epsilon}}_{\theta}$ is a function implemented by trainable models $\theta$ which can be seen as a noise predictor. A simplified version of objective function for a data point ${\bm{x}}$ used to train DDPMs is:
@@ 2.2 Data Attribution
Given a DDPM trained on dataset ${\mathbb{S}}$, our objective is to trace the influence of the training data on an output generated for a particular sample ${\bm{z}}$. This task is commonly referred to as data attribution. A common approach to data attribution involves addressing a counterfactual question: if a training sample ${\bm{z}}^{(i)}$ is removed from ${\mathbb{S}}$ and a model $\theta_{\backslash i}$ is retrained on the remaining subset ${\mathbb{S}}_{\backslash i}$, the influence of ${\bm{z}}^{(i)}$ on the sample ${\bm{z}}$ can be assessed by the change in the model output, computed as $f({\bm{z}},\theta)-f({\bm{z}},\theta_{\backslash i})$. The function $f({\bm{z}},\theta)$, which represents the model output, has a variety of choices, such as the direct output of the model or its loss function.
To circumvent the high computational costs associated with retraining models, some data attribution methods compute a scoring function $\tau({\bm{z}},{\mathbb{S}}):\mathcal{Z}\times\mathcal{Z}^{n}\rightarrow\mathbb{R}^{n}$, which assigns scores reflecting the importance of each training sample in ${\mathbb{S}}$ for the sample ${\bm{z}}$ under consideration. For clarity, $\tau({\bm{z}},{\mathbb{S}})^{(i)}$ denotes the attribution score assigned to the influence of the individual training sample ${\bm{z}}^{(i)}$ on ${\bm{z}}$.
TRAK (Park et al., 2023) stands out as a representative data attribution method designed for large-scale models focused on discriminative tasks. TRAK defines the model output function as:
@@
where $\hat{p}({\bm{x}},\theta)$ represents the probability of the corresponding class $y$ for the sample ${\bm{z}}$. TRAK introduces an attribution function $\tau_{\text{TRAK}}({\bm{z}},{\mathbb{S}})^{(i)}$ designed to approximate the change in $f_{\text{TRAK}}({\bm{z}},\theta)$ following the deletion of a data point. This is expressed as:
@@
where $r^{(i)}=1-\hat{p}({\bm{x}}^{(i)},\theta)$ denotes the residual for sample ${\bm{z}}^{(i)}$. Here, $\phi({\bm{z}}):={\bm{P}}^{\top}\nabla_{\theta}f_{\text{TRAK}}({\bm{z}},\theta)$ and ${\bm{\Phi}}:=[\phi({\bm{z}}^{(1)}),...,\phi({\bm{z}}^{(n)})]$ represents the matrix of stacked gradients from ${\mathbb{S}}$. ${\bm{P}}\sim\mathcal{N}(0,1)^{d\times k}$ is a random projection matrix (Johnson & Lindenstrauss, 1984) employed to reduce the dimension of the gradient $\nabla_{\theta}f_{\text{TRAK}}({\bm{z}};\theta)\in\mathbb{R}^{d}$ to $k$. This function computes a score indicating the influence of a training sample ${\bm{z}}_{i}$ on the sample of interest ${\bm{z}}$ in a discriminative model setting.
To adapt TRAK in diffusion models, D-TRAK (Zheng et al., 2024) modifies the output function to align with the objective function described in Eq. 3: $f_{\text{D-TRAK}}({\bm{z}},\theta)=\mathcal{L}_{\text{Simple}}({\bm{x}},\theta)$, and simplifies the residual term to an identity matrix ${\bm{I}}$. The attribution function in D-TRAK is defined as follows:
@@
where $\phi({\bm{z}}):={\bm{P}}^{\top}\nabla_{\theta}f_{\text{D-TRAK}}({\bm{z}},\theta)$ mirrors the gradient term used in TRAK. Interestingly, D-TRAK has observed that substituting the function $f_{\text{D-TRAK}}$ with other functions can yield superior attribution performance. Examples of these include $\mathcal{L}_{\text{Square}}({\bm{z}},\theta)=\mathbb{E}_{t,{\bm{\epsilon}}}[||{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)||^{2}]$ and $\mathcal{L}{\text{Average}}({\bm{z}},\theta)=\mathbb{E}_{t,{\bm{\epsilon}}}[\text{Avg}({\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t))]$, both of which have enhanced attribution capabilities.

### v4

@@ Abstract
As diffusion models become increasingly popular, the misuse of copyrighted and private images has emerged as a major concern. One promising solution to mitigate this issue is identifying the contribution of specific training samples in generative models, a process known as data attribution. Existing data attribution methods for diffusion models typically quantify the contribution of a training sample by evaluating the change in diffusion loss when the sample is included or excluded from the training process. However, we argue that the direct usage of diffusion loss cannot represent such a contribution accurately due to the calculation of diffusion loss. Specifically, these approaches measure the divergence between predicted and ground truth distributions, which leads to an indirect comparison between the predicted distributions and cannot represent the variances between model behaviors. To address these issues, we aim to measure the direct comparison between predicted distributions with an attribution score to analyse the training sample importance, which is achieved by Diffusion Attribution Score (DAS). Underpinned by rigorous theoretical analysis, we elucidate the effectiveness of DAS. Additionally, we explore strategies to accelerate DAS calculations, facilitating its application to large-scale diffusion models. Our extensive experiments across various datasets and diffusion models demonstrate that DAS significantly surpasses previous benchmarks in terms of the linear data-modelling score, establishing new state-of-the-art performance.  [^1] Code is available at https://github.com/Jinxu-Lin/DAS.
@@ 1 Introduction
Diffusion models, highlighted in key studies (Ho et al., 2020; Song et al., 2021b), are advancing significantly in generative machine learning with broad applications from image generation to artistic creation (Saharia et al., 2022; Hertz et al., 2023; Li et al., 2022; Ho et al., 2022). As these models, exemplified by projects like Stable Diffusion (Rombach et al., 2022), become increasingly capable of producing high-quality, varied outputs, the misuse of copyrighted and private images has become a significant concern. A key strategy to address this issue is identifying the contributions of training samples in generative models by evaluating their influence on the generation, a task known as data attribution. Data attribution in machine learning is to trace model outputs back to influential training examples, which is essential to understand how specific data points affect model behaviors. In practical applications, data attribution spans various domains, including explaining predictions (Koh & Liang, 2017; Yeh et al., 2018; Ilyas et al., 2022), curating datasets (Khanna et al., 2019; Jia et al., 2021; Liu et al., 2021), and dissecting the mechanisms of generative models like GANs and VAEs (Kong & Chaudhuri, 2021; Terashita et al., 2021), serving to enhance model transparency and explore the effect of training data on model behaviors.
@@
Data attribution methods generally fall into two categories. The first, based on sampling (Shapley et al., 1953; Ghorbani & Zou, 2019; Ilyas et al., 2022), involves retraining models to assess how outputs change with the a data deletion. While effective, this method requires training thousands of models. The second approach uses approximations to assess the change in output function for efficiency (Koh & Liang, 2017; Feldman & Zhang, 2020; Pruthi et al., 2020) by proposing attribution score function. TRAK (Park et al., 2023) introduced a innovative estimator, which considers the inverse-sigmoid function as model output and computes attribution scores to assess its change. Building on this, Zheng et al. (2024) proposed D-TRAK, adapting TRAK to diffusion models by following TRAK’s derivation and replacing the inverse-sigmoid function with the diffusion loss. Moreover, D-TRAK reported counterintuitive findings that the output function could be replaced with alternative functions without altering the score’s form. These empirical designs outperformed theoretically motivated diffusion losses in experiments, emphasizing the need for a deeper understanding of the attribution properties on diffusion models.
In this paper, we define the objective of data attribution in diffusion models as evaluating the impact of training samples on generation by measuring shifts in the predicted distribution after removing specific samples and retraining the model. From this perspective, directly applying TRAK to diffusion models by replacing the output function with the diffusion loss leads to indirect comparisons between predicted distributions, as the diffusion loss represents the KL-divergence between predicted and ground truth distributions (Ho et al., 2020). We also address the counter-intuitive findings in D-TRAK, which manually removes the effect of data distributions in attribution but treats the diffusion model’s output as a scalar, failing to capture the unique characteristics of diffusion models. Our analysis indicates that TRAK’s derivation cannot be directly extended to diffusion models, as they fundamentally differ from discriminative models. To address this, we propose the Diffusion Attribution Score (DAS), a novel metric designed specifically for diffusion models to quantify the impact of training samples by measuring the KL-divergence between predicted distributions when a sample is included or excluded from the training set. DAS computes this divergence through changes in the noise predictor’s output, which, by linearization, can be represented as variations in the model parameters. These parameter changes are then measured using Newton’s Method. Since directly computing the full form of DAS is computationally expensive, we propose several techniques, such as compressing models and datasets, to accelerate computation. These enhancements enable the application of DAS to large-scale diffusion models, significantly improving its practicality. Our experiments across ranges of datasets and diffusion models demonstrate that DAS significantly outperforms previous benchmarks, including D-TRAK and TRAK, achieving superior results in terms of the linear data-modeling score. This consistent performance across diverse settings highlights its robustness and establishes DAS as the new state-of-the-art method for data attribution in diffusion models. The primary contributions of our work are summarized as follows:
@@
    We provide a comprehensive analysis of the limitations of directly applying TRAK to diffusion models and evaluate D-TRAK’s empirical design, highlighting the need for more effective attribution methods tailored to diffusion models.
@@
2 Related Works
2.1 Data Attribution
The training data exerts a significant influence on the behavior of machine learning models. Data attribution aims to accurately assess the importance of each piece of training data in relation to the desired model outputs. However, methods of data attribution often face the challenge of balancing computational efficiency with accuracy. Sampling-based approaches, such as empirical influence functions (Feldman & Zhang, 2020), Shapley value estimators (Ghorbani & Zou, 2019; Jia et al., 2019), and datamodels (Ilyas et al., 2022), are able to precisely attribute influences to training data but typically necessitate the training of thousands of models to yield dependable results. On the other hand, methods like influence approximation (Koh & Liang, 2017; Schioppa et al., 2022) and gradient agreement scoring (Pruthi et al., 2020) provide computational benefits but may falter in terms of reliability in non-convex settings (Basu et al., 2021; Akyurek et al., 2022).
2.2 Data Attribution in Generative Models
The discussed methods address counterfactual questions within the context of discriminative models, focusing primarily on accuracy and model predictions. Extending these methodologies to generative models presents complexities due to the lack of clear labels or definitive ground truth. Research in this area includes efforts to compute influence within Generative Adversarial Networks (GANs) (Terashita et al., 2021) and Variational Autoencoders (VAEs) (Kong & Chaudhuri, 2021). In the realm of diffusion models, earlier research (Dai & Gifford, 2023) has explored influence computation by employing ensembles that necessitate training multiple models on varied subsets of training data—a method less suited for traditionally trained models.  Wang et al. (2023) suggest an alternative termed ”customization,” which involves adapting or tuning a pretrained text-to-image model through a specially designed training procedure. MONTAGE (Brokman et al., 2025) integrates a novel technique to monitor generations throughout the training via internal model representations. In this paper, we mainly focus on post-hoc data attribution method, which entails applying attribution methods after training. Recently, Park et al. (2023) developed TRAK, a new attribution method that is both effective and computationally feasible for large-scale models. Journey TRAK (Georgiev et al., 2023) extends TRAK to diffusion models by attributing influence across individual denoising timesteps. Moreover, D-TRAK (Zheng et al., 2024) has revealed surprising results, indicating that theoretically dubious choices in the design of TRAK might enhance performance, highlighting the imperative for further exploration into data attribution within diffusion models. In DataInf (Kwon et al., 2024), influence function using the loss gradient and Hessian have been improved for greater accuracy and efficiency in attributing diffusion models. A concurrent work (Mlodozeniec et al., 2025) extends influence functions to diffusion models by predicting how the probability of generating specific examples changes when training data is removed, leveraging a K-FAC approximation for scalable computation. These studies are pivotal in advancing our understanding and fostering the development of instance-based interpretations in unsupervised learning contexts. We give a more detailed theoretical discussion about the existing data attribution methods in diffusion model in Appendix E.3. Besides, we provide an introduction about application of data attribution in Appendix I.
@@ 3.1 Diffusion Models
Our study concentrates on discrete-time diffusion models, specifically Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) and Latent Diffusion Models (LDMs) which are foundational to Stable Diffusion (Rombach et al., 2022). This paper grounds all theoretical derivations within the framework of unconditional generation using DDPMs. Below, we detail the notation employed in DDPMs that underpins all further theoretical discussions.
Consider a training set ${\mathbb{S}}=\{{\bm{z}}^{(1)},...,{\bm{z}}^{(n)}\}$ where each training sample ${\bm{z}}^{(i)}:=({\bm{x}}^{(i)},y^{(i)})\sim\mathcal{Z}$ is an input-label pair [^2] In text-to-image task, ${\bm{z}}:=({\bm{x}}^{i},y^{i})$ represents an image-caption sample, whereas in unconditional generation, it solely contains an image ${\bm{z}}:=({\bm{x}})$.. Given an input data distribution $q({\bm{x}})$, DDPMs aim to model a distribution $p_{\theta}({\bm{x}})$ to approximate $q({\bm{x}})$. The learning process is divided into forward and reverse process, conducted over a series of timesteps in the latent variable space, with ${\bm{x}}_{0}$ denoting the initial image and ${\bm{x}}_{t}$ the latent variables at timestep $t\in[1,T]$. In the forward process, DDPMs sample an observation ${\bm{x}}_{0}$ from ${\mathbb{S}}$ and add noise on it across $T$ timesteps: $q({\bm{x}}_{t}|{\bm{x}}_{t-1}):=\mathcal{N}({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},\beta_{t}{\bm{I}})$, where $\beta_{1},...,\beta_{T}$ constitute a variance schedule. As indicated in DDPMs, the latent variable ${\bm{x}}_{t}$ can be express as a linear combination of ${\bm{x}}_{0}$:
@@
where $\alpha_{t}:=1-\beta_{t},\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$ and ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$. In the reward process, DDPMs model a distribution $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ by minimizing the KL-divergence from data at $t$:
@@
where ${\bm{\epsilon}}_{\theta}$ is a function implemented by models $\theta$ which can be seen as a noise predictor. A simplified version of objective function for a data point ${\bm{x}}$ used to train DDPMs is:
@@ 3.2 Data Attribution
Given a DDPM trained on dataset ${\mathbb{S}}$, our objective is to trace the influence of the training data on the generation of sample ${\bm{z}}$. This task is referred to as data attribution, which is commonly solved by addressing a counterfactual question: removing a training sample ${\bm{z}}^{(i)}$ from ${\mathbb{S}}$ and retraining a model $\theta_{\backslash i}$ on the subset ${\mathbb{S}}_{\backslash i}$, the influence of ${\bm{z}}^{(i)}$ on ${\bm{z}}$ can be assessed by the change in the model output, computed as $f({\bm{z}},\theta)-f({\bm{z}},\theta_{\backslash i})$. The function $f({\bm{z}},\theta)$, which represents the model output, has a variety of choices, such as the direct output of the model or the loss function.
To avoid the high costs of model retraining, some data attribution methods compute a score function $\tau({\bm{z}},{\mathbb{S}}):\mathcal{Z}\times\mathcal{Z}^{n}\rightarrow\mathbb{R}^{n}$ to reflect the importance of each training sample in ${\mathbb{S}}$ on the sample ${\bm{z}}$. For clarity, $\tau({\bm{z}},{\mathbb{S}})^{(i)}$ denotes the attribution score assigned to the individual training sample ${\bm{z}}^{(i)}$ on ${\bm{z}}$. TRAK (Park et al., 2023) stands out as a representative data attribution method designed for large-scale models focused on discriminative tasks. TRAK defines the model output function as:
@@
where $\hat{p}({\bm{x}},\theta)$ is the corresponding class probability of ${\bm{z}}$. TRAK then introduces attribution score $\tau_{\text{TRAK}}({\bm{z}},{\mathbb{S}})^{(i)}$ to approximate the change in $f_{\text{TRAK}}$ after a data intervention, which is expressed as:
@@
where $r^{(i)}:=[1-\hat{p}({\bm{x}}^{(i)},\theta)]$ denotes the residual for sample ${\bm{z}}^{(i)}$. Here, $\phi({\bm{z}}):={\bm{P}}^{\top}\nabla_{\theta}f_{\text{TRAK}}({\bm{z}},\theta)$ and ${\bm{\Phi}}:=[\phi({\bm{z}}^{(1)}),...,\phi({\bm{z}}^{(n)})]$ represents the matrix of stacked gradients on ${\mathbb{S}}$. ${\bm{P}}\sim\mathcal{N}(0,1)^{d\times k}$ is a random projection matrix (Johnson & Lindenstrauss, 1984) to reduce the gradient dimension.
To adapt TRAK in diffusion models, D-TRAK (Zheng et al., 2024) first followed TRAK’s guidance, replacing the output function with Simple Loss: $f_{\text{D-TRAK}}({\bm{z}},\theta)=\mathcal{L}_{\text{Simple}}({\bm{x}},\theta)$, and simplifies the residual term to an identity matrix ${\bm{I}}$. The attribution function in D-TRAK is defined as follows:
@@
where $\phi({\bm{z}}):={\bm{P}}^{\top}\nabla_{\theta}f_{\text{D-TRAK}}({\bm{z}},\theta)$. Interestingly, D-TRAK observed that substituting Simple Loss with other functions can yield superior attribution performance. Examples of these include $\mathcal{L}_{\text{Square}}({\bm{z}},\theta)=\mathbb{E}_{t,{\bm{\epsilon}}}[||{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)||^{2}]$ and $\mathcal{L}{\text{Average}}({\bm{z}},\theta)=\mathbb{E}_{t,{\bm{\epsilon}}}[\text{Avg}({\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t))]$.